A steel wire of length 64 cm weighs 5 g. If it is stretched by a force of 8 N, what would be the speed of a transverse wave passing on it?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. . mass = 5 g, length l = 64 cm ∴ mass per unit length = m = 5/64 g/cm ∴ Tension, T = 8N = 8 * 105 dyne V = √(T /m) = √(8 * 105 * 64) / 5 = 3200 cm/s = 32 m/s